OpenStudy (anonymous):

Hello! If anyone can help me to solve this problem, I would really appreaciate it! The problem: \frac{4}{5} \sqrt[3]{16}+ \frac{2}{25} \sqrt[3]{250}-4 \sqrt[3]{2}+ \sqrt[3]{2000} Thank you very much!

OpenStudy (anonymous):

\[\frac{4}{5} \sqrt[3]{16}+ \frac{2}{25} \sqrt[3]{250}-4 \sqrt[3]{2}+ \sqrt[3]{2000}\]